Guiana dolphins are one of the few mammals able to detect electric fields. In a test of sensitivity, a dolphin was exposed to the variable electric field from a pair of charged electrodes. The magnitude of the electric field near the sensory organs was measured by detecting the potential difference between two measurement electrodes located 1.0 cm apart along the field lines. The dolphin could reliably detect a field that produced a potential difference of 0.50 mV between these two electrodes. What is the corresponding electric field strength?
Answer #1

Given , distance (d) = 1 cm = 0.01 m , potential difference (V) = 0.5 mV = 0.5*10^-3 V

Using the formula,

E = V/d

E = 0.5*10^-3/0.01

E = 0.05 V/m

Thus , Electric field (E) is 0.05 V/m